# Espresso: Robust Concept Filtering in Text-to-Image Models

### v2

@@ 5.2. Concepts
Table 2. Summary of unacceptable concepts ($c^{u}$), acceptable concepts ($c^{a}$), and whether we use them to evaluate fine-tuning CRTs, filtering CRTs, or both.
@@
Identifying Acceptable Concepts. We specify a matching $c^{a}$ for every $c^{u}$ so that CRTs can steer away from generating $c^{u}$ concepts and instead focus on $c^{a}$. We indicate “$c^{u}$rightarrow $c^{a}$” for each group and discuss our choice of $c^{a}$.
  • Group-1: $c^{a}$ is the opposite of $c^{u}$, except for nudity. For instance, violence $\rightarrow$ peaceful, disturbing $\rightarrow$ pleasing, and hateful $\rightarrow$ loving. For nudity, we compared clothed with clean on a separate validation dataset (c.f. Section 5.3 and Table 4), and eliminated clothed because it blocked images with minimal exposed skin despite being acceptable (Qu et al., 2023b; Schramowski et al., 2023). Similarly, we evaluated and eliminated alternatives of $c^{a}$ for violence (nonviolent, and gentle), disturbing (calming, and soothing), and hateful (compassionate, and kind).
  • Group-2: $c^{a}$ is the type of $c^{u}$, taken from prior works (Kumari et al., 2023a; Heng and Soh, 2023b; Gandikota et al., 2023a). For instance, Grumpy Cat $\rightarrow$ cat, Nemo $\rightarrow$ fish, Captain Marvel $\rightarrow$ female superhero, Snoopy $\rightarrow$ dog, and R2D2 $\rightarrow$ robot.
[…]
Selecting Concepts to Evaluate CRTs. Different prior works used different concepts in their evaluation. Filtering CRTs were evaluated on inappropriate concepts in Group-1 while fine-tuning CRTs were evaluated on nudity and violence from Group-1, and all the concepts from Group-2 and Group-3. Hence, we compare Espresso with each CRT category separately using concepts they were evaluated on.
We summarize all the concepts used for evaluation in Table 2.
5.3. Dataset and Model Configurations
Training/Fine-tuning Datasets. We summarize the configuration for training/fine-tuning datasets in Table 3.
Table 3. Training dataset configurations required by CRTs.
CRT Configuration Requirements
CA (Kumari et al., 2023a) 200 unacceptable prompts ($\texttt{p}^{a}$) from ChatGPT with one image/prompt from SDv1.4
ESD (Gandikota et al., 2023a) Unacceptable concept ($c^{u}$)
FMN (Zhang et al., 2023b) 8 unacceptabe prompts ($\texttt{p}^{u}$) and one image/prompt from SDv1.4
SDD (Kim et al., 2023) Unacceptable concept ($c^{u}$) and 10 corresponding images
SA (Heng and Soh, 2023b) Acceptable concept ($c^{a}$) and 1000 corresponding images, and 6 unacceptable prompts ($\texttt{p}^{u}$)
UCE (Gandikota et al., 2024) Unacceptable and acceptable concepts ($c^{u}$ and $c^{a}$)
UD (Qu et al., 2023b) 776 total images with 580 acceptable images ($x^{a}$), and 196 unacceptable images ($x^{u}$): nudity (48), violence (45), disturbing (68), and hateful (35)
Fine-tuning CRTs. Different fine-tuning CRTs use different dataset configurations to ensure that R1 and R2 are met. We use the exact same configuration as described in each of the prior works to ensure that they satisfy the requirements claimed in their papers. While the configuration across different CRTs vary, the configuration for a specific CRT across different concepts is the same, in accordance with the instructions in the respective works. We briefly summarize them in Table 3. CA uses acceptable prompts which are generated from ChatGPT such that they contain $c^{a}$, and 1 image per prompt. ESD and SDD both use an unacceptable concept, and SDD additionally uses 10 images corresponding to $c^{u}$. All images are generated using SD v1.4. FMN uses unacceptable prompts of the form “An image of {$c^{u}$}”, and 1 image per prompt. For all the fine-tuning CRTs, we use their publicly available code for training to ensure the configuration is same as reported in their papers (Kumari et al., 2023b; Gandikota et al., 2023b; Zhang et al., 2023c; Gandikota et al., 2023c; Heng and Soh, 2023a; Kim et al., 2024; Qu et al., 2023a).
Filtering CRTs. We train $\texttt{F}_{\texttt{UD}}$ using its exact dataset configuration (Table 3) and code (Qu et al., 2023b). For Espresso, we follow the template by Kumari et al. (Kumari et al., 2023a) to generate 10 unacceptable and acceptable prompts using ChatGPT, with one image per prompt from SDv1.4. For Group-2 concepts, we randomly select 10 ChatGPT-generated prompts from the datasets provided by Kumari et al. (Kumari et al., 2023a). Our choice for having a small amount of training data is inspired from prior works on poisoning attacks which show that little data is required to modify CLIP (Yang et al., 2023; Carlini and Terzis, 2022). Our results also show that using such little data is indeed sufficient to meet R1, R2, and R3 (c.f. Section 6).
Validation Datasets. For R1, we generate 10 unacceptable prompts using ChatGPT, and generate 5 images per prompt using SD v1.4. We denote this validation dataset as $\mathcal{D}_{val}^{u}$. For R2, we randomly chose 100 non-overlapping acceptable prompts from the COCO 2014 dataset (Lin et al., 2014), and 1 image per prompt, generated from SD v1.4. We denote this validation dataset as $\mathcal{D}_{val}^{a}$. We summarize in Table 4.
Table 4. Configuration for Validation Datasets.
Concepts Dataset Configuration Requirements
Group-1 $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
Group-2 $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
Group-3 $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
All $\mathcal{D}_{val}^{a}$ 100 acceptable prompts ($\texttt{p}^{a}$) and 1 image ($x^{a}$) per prompt from the COCO 2014 dataset
Evaluation Datasets. We now describe the test datasets for evaluation: test dataset with acceptable prompts is $\mathcal{D}_{te}^{a}=\{(\texttt{$x^{a}$}_{j},\texttt{p}^{a}_{j})\}_{j=1}^{N}$ and with unacceptable prompts is $\mathcal{D}_{te}^{u}=\{(\texttt{$x^{u}$}_{j},\texttt{p}^{u}_{j})\}_{j=1}^{N}$.
For Group-1, we use an independent benchmark dataset, Inappropriate Image Prompts (I2P), containing prompts likely to generate unsafe images (Schramowski et al., 2023). It covers all four Group-1 concepts. For nudity, I2P includes a “nudity percentage” attribute. We choose all unacceptable prompts with a nudity percentage $>10$, resulting in a total of 300 unacceptable prompts, in accordance with prior works (Gandikota et al., 2023a, 2024). UD (Qu et al., 2023b) also used political as a concept, which is excluded from our evaluation since it is not a part of I2P. Note that I2P prompts do not explicitly contain the unacceptable concepts, i.e., the words nude, violent, hateful, or disturbing. For Group-2 and Group-3 concepts, there are no standard benchmark datasets. Hence, we generate the dataset by following prior works (Kumari et al., 2023a; Gandikota et al., 2023a) to obtain 200 unacceptable images, $x^{u}$, generated from SDv1.4 from 10 unacceptable prompts, $\texttt{p}^{u}$, generated from ChatGPT. These datasets constitute $\mathcal{D}_{te}^{u}$. For utility, we use the COCO 2014 test dataset ($\mathcal{D}_{te}^{a}$) with 200 randomly chosen acceptable prompts. This dataset constitutes $\mathcal{D}_{te}^{a}$. We summarize these datasets in Table 5.
Table 5. Configuration for Evaluation Datasets.
Concepts Dataset Configuration Requirements
Group-1 $\mathcal{D}_{te}^{u}$ I2P Dataset (Schramowski et al., 2023) of unacceptable prompts: for nudity: 449 prompts, violence: 758, disturbing: 857, hateful: 235
Group-2 $\mathcal{D}_{te}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 20 images ($x^{u}$) per prompt from SDv1.4
Group-3 $\mathcal{D}_{te}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 20 images ($x^{u}$) per prompt from SDv1.4
All $\mathcal{D}_{te}^{a}$ 200 acceptable prompts ($\texttt{p}^{a}$) from COCO 2014 dataset
@@ 5.3. Dataset and Model Configurations
Table 6. R1 Effectiveness: Comparison with fine-tuning CRTs using CLIP accuracy on unacceptable prompts (lower is better). We use red if accuracy is $>$50; blue if accuracy is between 25-50; green if accuracy is $<$25.
@@
Table 7. R2 (Utility): Comparison with fine-tuning CRTs using normalized CLIP scores on acceptable prompts (higher is better). We use red if score is between 50-70, blue if between 70-90; green if $>$90.
@@
Table 8. R3 (Robustness): Comparison with fine-tuning CRTs using CLIP accuracy on adversarial prompts (lower is better). We evaluate fine-tuning CRT’s against CCE and Espresso against CCE+ since CCE is already adapted to fine-tuning CRT’s. We use red if accuracy is $>$50; blue if accuracy is between 25-50; green if accuracy is $<$25.
@@
5.4. Metrics
We now describe the metrics to evaluate each of the requirements. We assume access to a reference CLIP, separate from those used in CRTs and f. This is a standard assumption in prior works to compute different metrics (Kumari et al., 2023a; Hessel et al., 2021; Kim et al., 2023; Gandikota et al., 2024).
R1 (Effectiveness). Depending on the CRT, we use:
  • CLIP accuracy (Hessel et al., 2021; Kumari et al., 2023a) for fine-tuning CRTs. This is the cosine similarity (divided by the temperature parameter $\tau$) between the embeddings of the generated image $x$ with the embeddings of $c^{u}$ and $c^{a}$ from the reference CLIP, where $\tau$ is the temperature parameter of the reference CLIP, $\frac{1}{100}$. This outputs the likelihood of predicting $c^{u}$. Hence, CLIP accuracy should be low (ideally zero) for effective concept removal. Formally, it is:
    $\frac{\text{exp}{(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{u})/\tau))}}{\text{exp}(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{u})/\tau))+\text{exp}(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{a})/\tau))}$
    where $\tilde{\phi_{p}}$ and $\tilde{\phi_{x}}$ are embeddings from the reference CLIP. If Espresso detects $x^{u}$, a replacement image is generated. We calculate the CLIP accuracy on the final set of images after filtering. Since the metrics compare with baseline reference CLIP, we use the same default replacement image as in the SD v1.4 filter.
  • False Negative Rates (FNR) (Qu et al., 2023b; Rando et al., 2022) for filtering CRTs is the fraction of images with $c^{u}$ which are not blocked. It should be low (ideally zero).
R2 (Utility). Depending on the CRT, we use:
  • Normalized CLIP score for fine-tuning CRTs, which is the ratio of cosine similarity (scaled by $\frac{1}{\tau}$, where $\tau=\frac{1}{100}$) between $\phi_{x}(\texttt{$x$})$ and $\phi_{p}(\texttt{p})$ from f compared to that from a reference CLIP as a baseline. Formally, $\frac{\text{cos}(\phi_{x}(\texttt{$x$}),\phi_{p}(\texttt{p}))/\tau}{\text{cos}(\tilde{\phi}_{x}({\texttt{$x$}})/,\tilde{\phi}_{p}(\texttt{p}))/\tau}$. Normalized CLIP score should be high (ideally one) for high utility. This metric is different from standard CLIP score from prior work (Hessel et al., 2021; Kumari et al., 2023a; Kim et al., 2023; Gandikota et al., 2024) which measures the cosine similarity between $\phi_{x}(\texttt{$x$})$ and $\phi_{p}(\texttt{p})$ from f. Normalizing the CLIP score helps compare f to the reference CLIP which has the maximum achievable CLIP score.
  • False Positive Rates (FPR) (Qu et al., 2023b; Rando et al., 2022) for filtering CRTs is the fraction of images without $c^{u}$ which are blocked. It should be low (ideally zero).
R3 (Robustness). We use the same metrics as effectiveness: CLIP accuracy to evaluate fine-tuning CRTs and compare Espresso with them, and FNR for filtering CRTs. However, we compute both the metrics on $\mathcal{D}_{adv}^{u}$.
6. Evaluation
@@
We now evaluate Espresso, first comparing with fine-tuning CRTs (Section 6.1), followed by filtering CRT (Section 6.2).
6.1. Comparison with Fine-tuning CRTs
R1 Effectiveness. We report CLIP accuracy in Table 6 for all eleven concepts on $\mathcal{D}_{te}^{u}$.
@@ 6.1. Comparison with Fine-tuning CRTs
All CRTs show poor accuracy on nudity and violence. We attribute this to fine-tuning CRTs being sensitive to the input prompts (Pham et al., 2024b; Yang et al., 2024a; Ma et al., 2024). Hence, they perform poorly for nudity and violence which do not explicitly contain unacceptable concepts. In comparison to other CRTs, Espresso consistently maintains high accuracy on the I2P benchmark dataset as it classified based on the generated images and not prompts. Further, Espresso satisfies R1 using only 20 prompts and corresponding images for training. This is consistent with results on CLIP poisoning where changing the performance of CLIP only required a small number of poison samples (Carlini and Terzis, 2022; Yang et al., 2023).
@@
ESD, UCE, and SDD have better accuracy compared to the other three fine-tuning CRTs. This could be attributed to the similar optimizations used by ESD and SDD for fine-tuning the T2I model. Both ESD and SDD fine-tune $\epsilon_{\theta}$, conditioned on $c^{u}$, to match original $\epsilon_{\theta}$ without any prompt to reduce the influence of $c^{u}$ on the output. For UCE, the reason for its higher effectiveness might be because it directly removes the influence of $c^{u}$ from the weights of T2I.
R2 Utility. We report normalized CLIP scores in Table 7 on $\mathcal{D}_{te}^{a}$.
@@
All the fine-tuning CRTs perform well across all concepts (either blue or green). This could be attributed to fine-tuning while explicitly including utility preservation as a requirement. We observe that CA with KL-Divergence-based optimization targeted for cross-attention layers, and UCE with a precise closed-form solution to model updates, preserve utility better than other CRTs.
Espresso achieves high utility preservation for all concepts except for violence, and some Group-2 concepts (Nemo, Captain Marvel, Snoopy, and R2D2). During fine-tuning, effectiveness improves but at the cost of utility, which is inherent to all fine-tuning CRTs. For violence, in our experiments, we observed an early decrease in utility during the very first epoch resulting in poor trade-off between effectiveness and utility preservation. We attribute the poor utility preservation on Group-2 concepts to the ambiguity in the unacceptable concepts. For instance, Nemo is both a fish and a ship captain (VERNE, 2024), and Captain Marvel represents both a male and a female superhero (Friedwald, 2019). To verify this, we precisely specify the unacceptable concepts to reduce ambiguity: as Nemo fish, Captain Marvel female superhero, Snoopy dog, and R2D2 robot. We evaluate Espresso on a separate validation dataset and then report the numbers for the evaluation dataset below: compared to the results in Table 7, the normalized CLIP score for this new configuration is: 0.97 $\pm$ 0.00 (Nemo), 0.90 $\pm$ 0.02 (Captain Marvel), 0.98 $\pm$ 0.03 (Snoopy), 0.92 $\pm$ 0.02 (R2D2), which are now indicated in green. We also report the CLIP accuracy to evaluate effectiveness with this new configuration: 0.02 $\pm$ 0.00 (Nemo), 0.00 $\pm$ 0.00 (Captain Marvel), 0.02 $\pm$ 0.01 (Snoopy), 0.00 $\pm$ 0.00 (R2D2), which are effective, same as before. Hence, precisely specifying $c^{u}$ is important for Espresso and can improve utility preservation without sacrificing effectiveness. In summary, Espresso preserves utility on most concepts, and the poor utility can be improved by a careful choice of $c^{u}$.
@@
R3 Robustness. We report CLIP accuracy in Table 8 on $\mathcal{D}_{adv}^{u}$. We evaluate different CRTs against four attacks: Typo+, PEZ+, CCE/CCE+, and RingBell+. We evaluate fine-tuning CRT’s against CCE, and evaluate Espresso against CCE+ as CCE already accounts for fine-tuning CRTs. We use the same color coding as in Table 6.
@@
CCE/CCE+, being a white-box attack which uses the parameters of the entire T2I model, is the strongest attack and it makes all fine-tuning CRTs ineffective. However, Espresso is effective against CCE/CCE+ and outperforms all fine-tuning CRTs. On the remaining attacks, all the fine-tuning CRTs have better robustness on Group-3 concepts than on Group-1 and 2. We attribute this to the difficulty of T2I models in generating precise faces while also evading detection, as shown in prior works (Matthias, 2023). Similar to results in Table 6, we note that all CRTs perform poorly on nudity and violence across all the attacks. This is expected as the CRTs have poor effectiveness on these concepts as seen in Table 6. Hence, adversarial prompts for these concepts are likely to easily evade them. In summary, Espresso is more robust than all prior fine-tuning CRTs on all the concepts across all the attacks. We discuss the possibility of future attacks against Espresso in Section 9.2.
@@
Summary. We depict the trade-offs among R1, R2, and R3 in Figure 1. We use (1-CLIP accuracy) for R1 on $\mathcal{D}_{te}^{u}$, normalized CLIP score on $\mathcal{D}_{te}^{a}$ for R2, and (1-CLIP accuracy) for R3 on $\mathcal{D}_{adv}^{u}$ using CCE/CCE+. For each of the requirements, we use the average across all concepts as the representative value for a CRT. Overall, Espresso provides a better trade-off across the three requirements compared to all the other CRTs.
@@
Figure 1. Espresso has a better trade-off across R1, R2, and R3 compared to other fine-tuning CRTs.
6.2. Comparison with Filtering CRT

### v4

@@ 5.4 Pipeline for Evaluating CRTs
TABLE IV: Summary of unacceptable concepts ($c^{u}$), acceptable concepts ($c^{a}$), and whether we use them to evaluate fine-tuning CRTs, filtering CRTs, or both.
@@
TABLE V: R1 Effectiveness: Comparison with fine-tuning CRTs using CLIP accuracy on unacceptable prompts (lower is better). We use red if accuracy is $>$50; blue if accuracy is between 25-50; green if accuracy is $<$25.
@@
TABLE VI: R2 (Utility): Comparison with fine-tuning CRTs using normalized CLIP scores on acceptable prompts (higher is better). We use red if score is between 50-70, blue if between 70-90; green if $>$90.
@@
For concepts in Group-1, we evaluate the effectiveness (R1) and utility (R2) for different candidates of $c^{a}$ on $\mathcal{D}_{val}^{u}$ and $\mathcal{D}_{val}^{a}$ respectively. We found that the following concepts gave the best results: nudity $\rightarrow$ clean, violence $\rightarrow$ peaceful, disturbing $\rightarrow$ pleasing, and hateful $\rightarrow$ loving. For nudity, we eliminated clothed as it blocked images with minimal exposed skin despite being acceptable [67, 52].
We use concepts in Group-1 following prior work [67] to evaluate filtering CRTs. We use nudity and violence from Group-1, and all the concepts from Group-2 and Group-3 following prior works [30, 21, 68, 26] to evaluate fine-tuning CRTs. Hence, we compare Espresso with each CRT category separately using concepts they were evaluated on. We summarize the final concepts for evaluation in Table IV.
Step 5: Evaluating R1, R2, and R3 on $\mathcal{D}_{te}$
We evaluate effectiveness (R1) on $\mathcal{D}_{te}^{u}$, utility (R2 on $\mathcal{D}_{te}^{a}$, and robustness (R3) $\mathcal{D}_{adv}^{u}$ which is generated by running different attacks $\mathcal{D}_{te}^{u}$ to get $\texttt{p}^{adv}$. In all cases, we only pass the prompts from the dataset to T2I models and compute different metrics described in Section 5.2.
6 Evaluation
TABLE VII: R3 (Robustness): Comparison with fine-tuning CRTs using CLIP accuracy on adversarial prompts (lower is better). We evaluate fine-tuning CRT’s against CCE and Espresso against CCE+ since CCE is already adapted to fine-tuning CRT’s. We use red if accuracy is $>$50; blue if accuracy is between 25-50; green if accuracy is $<$25.
@@ 6 Evaluation
We evaluate Espresso with fine-tuning CRTs (Section 6.1), and then filtering CRT (Section 6.2).
6.1 Comparison with Fine-tuning CRTs
@@
R1 Effectiveness. We report CLIP accuracy in Table V for all eleven concepts on $\mathcal{D}_{te}^{u}$. All CRTs show poor accuracy on nudity and violence. We attribute this to fine-tuning CRTs being sensitive to the input prompts [43, 65, 37]. Hence, they perform poorly for nudity and violence which do not explicitly include $c^{u}$. In comparison to other CRTs, Espresso consistently maintains high accuracy on the I2P benchmark dataset as it classifies the generated images and does not depend on the prompts. Further, Espresso satisfies R1 using only 20 prompts and corresponding images for training. This is consistent with results on CLIP poisoning where changing the performance of CLIP only required a small number of poison samples [7, 64].
@@
ESD, UCE, and SDD have better accuracy compared to the other three fine-tuning CRTs. This could be attributed to the similar optimizations used by ESD and SDD for fine-tuning the T2I model: both fine-tune $\epsilon_{\theta}$, conditioned on $c^{u}$, to match original $\epsilon_{\theta}$ without any prompt to reduce the influence of $c^{u}$ on the output. For UCE, higher effectiveness can be attributed to directly removing the influence of $c^{u}$ from the T2I model parameters. Overall, Espresso is more effective than other fine-tuning CRTs.
@@
R2 Utility. We report normalized CLIP scores in Table VI on $\mathcal{D}_{te}^{a}$. All the fine-tuning CRTs perform well across all concepts (either blue or green). This could be attributed to explicitly fine-tuning for utility. We observe that CA with KL-Divergence-based optimization targeted for cross-attention layers, and UCE with a precise closed-form solution to model updates, preserve utility better than others.
[…]
@@
R3 Robustness. We report CLIP accuracy in Table VII on $\mathcal{D}_{adv}^{u}$. We evaluate different CRTs against four attacks: Typo+, PEZ+, CCE/CCE+, and RingBell+. We evaluate fine-tuning CRT’s against CCE, and evaluate Espresso against CCE+ as CCE already accounts for fine-tuning CRTs. We use the same color coding as in Table V.
@@
CCE/CCE+, being a white-box attack which uses the parameters of the entire T2I model, is the strongest attack and it makes all fine-tuning CRTs ineffective. However, Espresso is more robust against CCE/CCE+ and outperforms all fine-tuning CRTs. On the remaining attacks, all the fine-tuning CRTs have better robustness on Group-3 concepts than on Group-1 and 2. We attribute this to the difficulty of T2I models in generating precise faces while also evading detection, as shown in prior works [39]. Similar to results in Table V, we note that all CRTs perform poorly on nudity and violence across all the attacks. This is expected as the CRTs have poor effectiveness on these concepts as seen in Table V. Hence, adversarial prompts for these concepts are likely to easily evade them. Overall, Espresso is more robust than all prior fine-tuning CRTs on all the concepts across all the attacks. We discuss the possibility of future attacks against Espresso in Section 9.1.
@@
Summary. We depict the trade-offs among R1, R2, and R3 in Figure 2. We use (1-CLIP accuracy) for R1 on $\mathcal{D}_{te}^{u}$, normalized CLIP score on $\mathcal{D}_{te}^{a}$ for R2, and (1-CLIP accuracy) for R3 on $\mathcal{D}_{adv}^{u}$ using CCE/CCE+. For each of the requirements, we use the average across all concepts as the representative value for a CRT. Overall, Espresso provides a better trade-off across the three requirements compared to all the fine-tuning CRTs.
@@
Figure 2: Espresso has a better trade-off than other fine-tuning CRTs.
6.2 Comparison with Filtering CRT